4 Titration of 0.2121 g of pure Na2C2O4 required 43.31 mL of KMnO4. What is the molar concentration of the KMnO4 solution? The chemical reaction is MnO4- + C2O42- + H+ ----------> Mn2+ + CO2 + H2O (unbalanced equation) Atomic masses: Na = 23 C = 12 O = 1

1
Expert's answer
2022-05-20T12:39:05-0400

Solution:

Balanced chemical equation:

2MnO4 + 5C2O42− + 16H+ → 2Mn2+ + 10CO2 + 8H2O

According to stoichiometry:

Moles of MnO4 / 2 = Moles of C2O42− / 5

5 × Moles of KMnO4 = 2 × Moles of Na2C2O4

5 × Molarity of KMnO4 × Volume of KMnO4 = 2 × (Mass of Na2C2O4 / Molar mass of Na2C2O4)


Molarity of KMnO4 = unknown

Volume of KMnO4 = 43.31 mL × (1 L / 1000 mL) = 0.04331 L

Mass of Na2C2O4 = 0.2121 g

Molar mass of Na2C2O4 = 2×Ar(Na) + 2×Ar(C) + 4×Ar(O) = 2×23 + 2×12 + 4×16 = 134 (g mol−1)


Therefore,

Molarity of KMnO4 = (2 × Mass of Na2C2O4) / (5 × Volume of KMnO4 × Molar mass of Na2C2O4)

Molarity of KMnO4 = (2 × 0.2121 g) / (5 × 0.04331 L × 134 g mol−1) = 0.01462 mol L−1

Molarity of KMnO4 = 0.01462 M


Answer: The molar concentration of the KMnO4 solution is 0.01462 M